# Pathways on the Image Manifold: Image Editing via Video Generation

### v1

@@ Abstract
Recent advances in image editing, driven by image diffusion models, have shown remarkable progress. However, significant challenges remain, as these models often struggle to follow complex edit instructions accurately and frequently compromise fidelity by altering key elements of the original image. Simultaneously, video generation has made remarkable strides, with models that effectively function as consistent and continuous world simulators. In this paper, we propose merging these two fields by utilizing image-to-video models for image editing. We reformulate image editing as a temporal process, using pretrained video models to create smooth transitions from the original image to the desired edit. This approach traverses the image manifold continuously, ensuring consistent edits while preserving the original image’s key aspects. Our approach achieves state-of-the-art results on text-based image editing, demonstrating significant improvements in both edit accuracy and image preservation.
@@ 1 Introduction
In this paper, we propose a paradigm shift in image editing by reformulating it as a video generation task. Instead of attempting this single state transition, our approach harnesses temporal coherence: the source image serves as the initial frame of a video that progressively and naturally transforms toward the target edit. This temporal evolution allows the editing process to unfold through physically plausible intermediate states, providing a continuous path between source and target images, as illustrated in Figure 1. This temporal approach leverages the sophisticated world understanding embedded in recent video generation models, which have achieved breakthrough results in temporal coherence and visual quality through training on large-scale internet data [46, 2]. From a geometrical perspective, conventional editing approaches project initial noise onto the natural image manifold, targeting a single point where the image aligns with both the source and the edit request. In contrast, our approach generates a continuous path along the manifold between the original and edited image, producing a smooth realistic transition across different image states, as thoroughly discussed in Section 4.
@@
Figure 2: Editing Manifold Pathway. Given an input image and target caption ”A happy person making a heart shape with their hands”, our method generates a continuous path on the natural image manifold. Each generated frame (indicated by black arrows) represents a plausible intermediate state between the source and target, maintaining temporal consistency throughout the transformation. As a result, in contrast to the competing approach, F2F achieves the desired edit while preserving the ”AI” text on the person’s shirt.
@@ 2.1 Image Editing
Text-based image editing has advanced significantly with the success of generative diffusion models [18]. These models, in their text-to-image version, generate images through a denoising diffusion process conditioned on input text, relying on ground truth pairs of text and images for training [37, 35]. In contrast, image editing often lacks predefined ground truth data for source and target images, presenting unique challenges. This limitation has led researchers to explore a diverse range of editing methodologies [19]. For example, SDEdit [28] introduces noise into an image and then denoises it based on an editing target prompt. Imagic [21] fine-tunes a text-to-image model on a single image, subsequently interpolating between input and target text embeddings to produce edits. Other methods first invert the input image into a diffusion model’s latent space [30, 20] and then generate the edited image from that latent representation using various techniques for structure preservation and manipulation [17, 32]. InstructPix2Pix [5] synthesizes an editing dataset using the approach in [17], filters it, and employs this dataset to train a diffusion model in a supervised fashion. Paint-by-Inpaint [45] further explores this supervised approach, generating a real-image dataset for object insertion.
@@ 2.2 Generative Video Models
Recent years have seen remarkable advancements in video generation, progressing from systems limited to specific domains [15] and brief clips to models capable of generating diverse, high-fidelity content. This progress has been notably driven by paradigm shifts and the significant scaling of both datasets [7] and architectures. Approaches have evolved from recurrent networks [16, 40] and generative adversarial networks (GANs) [42, 41, 9] to latent diffusion models (LDMs) [3, 2, 11], which leverage large U-Net or transformer-based architectures alongside vast internet-sourced datasets. Notable recent efforts include Stable Video Diffusion [2], which utilizes large curated datasets to train a latent diffusion model, and OpenAI’s Sora [26], which demonstrates impressive capabilities by reportedly leveraging a large architecture and extensive public and private datasets. These models have been described as ”world simulators” due to their emergent understanding of physical dynamics and temporal coherence. Our work builds upon CogVideoX [46], a transformer-based latent diffusion model that employs a 3D Variational Autoencoder to compress videos across both spatial and temporal dimensions, enhancing coherence. To improve text-video alignment, CogVideoX also integrates an expert transformer with adaptive LayerNorm, allowing deep fusion between visual and textual modalities.
@@ 3.1 Temporal Editing Captions
To automate this process, we leverage recent advances in Vision-Language Models (VLMs) [10, 44, 43, 14, 36]. The VLM, specifically ChatGPT-4o [31], is instructed to produce a concise video scenario that highlights how elements within the image change or move over time. The generated caption captures the essential transformations while maintaining a static camera perspective unless movement is necessary. To improve generation quality, we utilize in-context learning (ICL) [12], providing the VLM with nine exemplar prompt-caption pairs. The complete prompt template, ICL examples, and an ablation study comparing this approach with directly using the target captions are included in the Appendix.
@@ 3.2 Video Generation
We employ CogVideoX [46], a pretrained generative video latent diffusion model utilizing a transformer-based architecture. Specifically, we use its image-to-video variant, which has been fine-tuned to generate videos starting from an input image $I_{s}$. During generation, $I_{s}$ is encoded and concatenated with noise in latent space, where the model applies a denoising process guided by the temporal caption $\tilde{c}$. As elaborated in Section 4, this conditioning allows generated videos to start from $I_{s}$ and evolve naturally along the image manifold, maintaining temporal coherence and consistency. Additionally, the model’s transformer architecture enables effective fusion of visual and textual information, allowing precise control over the editing process through our temporal captions. Formally, given the video generator $G$, we define the generation process as:
@@ 3.3 Frame Selection
To automate the selection of $t^{*}$ and avoid manual frame-by-frame review, we employ an automated approach. After generating the sequence $V$, we sample every fourth frame, imprinting each with a unique identifier and assembling them into an image collage alongside $I_{s}$. Inspired by [22], which introduces a novel approach to video comprehension by transforming videos into image grids, we use a VLM, specifically GPT-4o, to assist in selecting $t^{*}$ by providing it with the collage and the editing prompt $c$. The VLM is tasked with identifying the frame that best fulfills the editing intent, evaluating each frame’s alignment with $c$ and fidelity to $I_{s}$. The model is instructed to select the optimal frame with the lowest index that completes the edit. An ablation study, detailed in the appendix, evaluates the effectiveness of our automated approach against the baseline of selecting the final frame of the video.
It is important to note that while we refer to an optimal $t^{*}$, the definition of the required edit can be subjective and dependent on user preferences. For example, in Figure 3, different users may select frames of the cat jumping based on variations in its altitude. Thus, frame selection could serve as a flexible and customizable advantage in certain scenarios.
@@ 4 Editing Manifold Pathway
We manually filter the outcomes to ensure alignment with the descriptions, with examples provided in the supplementary material. Then, we use a CLIP ViT-B/32 model [34] to extract a 512-dimensional feature vector for each image. Using these features, along with 25 feature vectors extracted from random noise images, we perform Principal Component Analysis (PCA) to reduce the dimensionality into a two-dimensional subspace. The resulting subspace, now suitable for visualization, is depicted in Figure 2.
@@
Consider an original image from the AI group that we wish to edit using the target prompt: “A happy person making a heart shape with their hands”. Current editing methods generate a single image, which may result in an abrupt transition to the red cluster, effectively removing the ’AI’ on the shirt. This behavior is illustrated in the figure, where the edit performed by LEDITS++ is positioned near the red cluster, and the ’AI’ text on the shirt disappears. In contrast, our method leverages video generation to perform the edit smoothly, moving along the manifold with incremental changes until it reaches the required edit in the purple cluster, as depicted by the black arrows in Figure 2. The temporal editing caption guiding this process is: “A happy person very slowly raising their hands to form a heart shape”. This gradual progression enables us to reach the purple cluster, resulting in an edited image where the person maintains the ‘AI’ on their shirt while making a heart shape—a faithful preservation of the original image’s key attributes. This experiment illustrates that the proposed paradigm enables smooth traversal across the image manifold, enabling consistent edits while preserving the essential characteristics of the original image.
@@ 5 Experiments
We evaluate Frame2Frame against state-of-the-art image editing methods, including LEdits++ [4], SDEdit [29], Pix2Pix-Zero [33] and Imagic [21]. Our evaluation spans two benchmarks: the established TEdBench [21] for general image editing, and our newly introduced PosEdit benchmark, specifically designed to evaluate human pose editing tasks. Finally, we conduct a human evaluation to better understand our method’s performance based on real user preferences.
@@ 5.1 Evaluation Protocol
Table 2: PosEdit Results. Quantitive evaluation on PosEdit. Source metrics assess similarity to the original image, while Target metrics now include both LPIPS and CLIP-I comparisons to the ground-truth target image, along with the CLIP score for edit accuracy.
@@ 5.2 TEdBench Evaluation Results
The qualitative advantages of our method are visually evident in the comparisons presented in Figure 4. The figure demonstrates our method’s superior performance across diverse editing scenarios, producing results that both faithfully execute the intended edits while maintaining strong alignment with the source image. Additional examples showcasing these capabilities are provided in the appendix.
@@ 5.3 PosEdit Benchmark
The proposed benchmark specifically focuses on human pose manipulation. Unlike TEdBench, each editing task in PosEdit includes a ground truth frame extracted from the same subject in the target pose. The availability of reference images allows for a more comprehensive evaluation of both editing accuracy and identity preservation. In the dataset, each editing task consists of two images: a source image showing the subject in a neutral standing pose with arms relaxed at their sides, and a ground-truth edited target image capturing the subject performing in a specific pose. The benchmark also provides a prompt for each task that describes the target pose the subject should achieve (e.g., ”A person in a basketball shooting posture.”).
Figure 5: Qualitative Results on PosEdit. Comparison between our Frame2Frame method and LEDITS++ on human motion editing tasks. For each example, we show the source image, edited results from both methods, and the ground-truth target image. Our method better preserves subject identity while achieving more natural pose transitions.
@@
The evaluation on PosEdit follows the same protocol established for TEdBench: generating multiple variants using nine different seeds and manually selecting the best result for each method. However, PosEdit’s ground-truth target images enable additional evaluation metrics. Beyond measuring preservation of source content (via LPIPS and CLIP-I between source and edited images) and edit accuracy (via CLIP score with the prompt), we compute LPIPS and CLIP-I metrics between the edited image and its corresponding ground-truth target. This enhanced evaluation directly assesses both the accuracy of the transformation and preservation of identity features.
@@
Two important implementation notes: First, for Pix2Pix-Zero, which requires source image descriptions, we use the same static prompt across all tasks: ”A person standing naturally with his arms relaxed at his sides.” Second, we exclude Imagic [21] from this comparison as their official implementation is not publicly available and their results are only reported on TEdBench.
@@
Table 3: Human Survey Results. Human evaluation on TEdBench shows that F2F surpasses LEDITS++ in edit accuracy while demonstrating a significant advantage in preserving the original image content.
5.4 Human Evaluation Survey
To further evaluate our method, we conducted a human survey. As indicated in Table 1, LEDITS++ emerges as the most competitive method compared to ours, making it a natural choice for comparison. Participants were presented with 20 random TEdBench samples, including the source image, the target edit prompt, and the corresponding edited outputs from both methods. Inspired by the methodology proposed by [47], participants evaluated each comparison based on two criteria: (1) edit accuracy relative to the prompt, and (2) edit quality—defined as the preservation of visual fidelity to the source image, seamless integration of edited elements, and the overall natural appearance of modifications. We collected responses from 59 randomly selected online evaluators. To ensure an unbiased assessment, evaluators were not informed of the study’s objectives. The results were quantified using two metrics: (i) overall global preference, expressed as a percentage, and (ii) aggregated per-image preference, where a tie resulted in each method receiving 0.5 points.
Our results, summarized in Table 3, demonstrate that F2F outperforms LEdits++ on both metrics. Specifically: For edit quality, F2F achieved a global preference score of 53%, slightly surpassing the 47% obtained by LEdits++. In terms of edit accuracy, F2F achieved a 65.6% global preference compared to LEdits++’s 34.4%. The per-image preference results mirrored this trend, indicating robustness across individual examples and no significant influence of outliers. These findings reinforce our claim that smooth temporal editing—using video as a medium—preserves essential scene characteristics while successfully performing the edit. Our results suggest that the gap between F2F and LEdits++ in source preservation is larger than reflected by the LPIPS scores in Table 1. The full survey details are provided in the supplementary materials.
@@ 5.5 Additional Vision Tasks
We demonstrate our framework’s applicability beyond traditional editing by applying it to fundamental image manipulation tasks: denoising, deblurring, outpainting, and relighting. For these experiments, we utilize Runway Gen-3 [13] as our video generation backbone, as it showed superior performance on these specific tasks compared to other generative video models. Our results, shown in Figure 6, demonstrate strong performance across all tasks. We associate this success to the natural alignment between these operations and common video scenarios: deblurring maps to camera focusing, outpainting to camera motion, and relighting to time-lapse lighting changes. This alignment enables our method to leverage the video model’s learned temporal dynamics, achieving high-quality results without task-specific training. Full prompts are provided in the supplementary material.

### v4

@@ Abstract
Recent advances in image editing, driven by image diffusion models, have shown remarkable progress. However, significant challenges remain, as these models often struggle to follow complex edit instructions accurately and frequently compromise fidelity by altering key elements of the original image. Simultaneously, video generation has made remarkable strides, with models that effectively function as consistent and continuous world simulators. In this paper, we propose merging these two fields by utilizing image-to-video models for image editing. We reformulate image editing as a temporal process, using pretrained video models to create smooth transitions from the original image to the desired edit. This approach traverses the image manifold continuously, ensuring consistent edits while preserving the original image’s key aspects. Our approach achieves state-of-the-art results on text-based image editing, demonstrating significant improvements in both edit accuracy and image preservation. Visit our project page.
@@ 1 Introduction
In this paper, we propose a paradigm shift in image editing by reformulating it as a video generation task. Rather than a single-state transition, our approach harnesses temporal coherence: the source image serves as the initial frame of a video that progressively and naturally transforms toward the target edit. This temporal evolution allows the editing process to unfold through physically plausible intermediate states, providing a continuous path between source and target images, as illustrated in Figure 1 and Figure S9 in the appendix. This temporal approach leverages the sophisticated world understanding embedded in recent video generation models, which have achieved breakthrough results in temporal coherence and visual quality through training on large-scale internet data [48, 2]. From a geometrical perspective, conventional editing approaches project initial noise onto the natural image manifold, targeting a single point where the image aligns with both the source and the edit request. In contrast, our approach generates a continuous path along the manifold between the original and edited image, producing a smooth realistic transition across different image states, as thoroughly discussed in Section 4.
@@
Figure 2: Editing Manifold Pathway. Given an input image and target caption "A happy person making a heart shape with their hands", our method generates a continuous path on the natural image manifold. Each generated frame (indicated by black arrows) represents a plausible intermediate state between the source and target, maintaining temporal consistency throughout the transformation. As a result, in contrast to the competing approach, F2F achieves the desired edit while preserving the "AI" text on the person’s shirt.
@@ 2.1 Image Editing
Text-based image editing has advanced significantly with the success of generative diffusion models [19]. These models, in their text-to-image version, generate images through a denoising diffusion process conditioned on input text, relying on ground truth text-image pairs for training [39, 37]. In contrast, image editing often lacks predefined ground truth data for source and target images, posing unique challenges. This limitation has led researchers to explore diverse editing methodologies [20]. For example, SDEdit [30] injects noise into an image and then denoises it based on an editing target prompt. Imagic [22] fine-tunes a text-to-image model on a single image, subsequently interpolating between input and target text embeddings to produce edits. Other methods first invert the input image into a diffusion model’s latent space [32, 21] and then generate the edited image from that latent representation using various techniques for structure preservation and manipulation [18, 34]. InstructPix2Pix [5] synthesizes an editing dataset using the approach in [18], filters it, and employs this dataset to train a diffusion model in a supervised fashion. Paint-by-Inpaint [47] further explores this supervised approach, generating a real-image dataset for object insertion.
@@ 2.2 Generative Video Models
Recent years have seen remarkable advancements in video generation, evolving from domain-specific systems [15] and brief clips to models capable of generating diverse, high-fidelity content. This progress has stemmed from paradigm shifts and the large-scale expansion of datasets [7] and architectures. Approaches have evolved from recurrent networks [16, 42] and generative adversarial networks (GANs) [44, 43, 9] to latent diffusion models (LDMs) [3, 2, 11], which leverage large U-Net or transformer-based architectures alongside vast internet-sourced datasets. Notable recent efforts include Stable Video Diffusion [2], which trains an LDM on large curated datasets, and OpenAI’s Sora [28], which achieves impressive results by integrating large architectures with extensive public and private datasets. These models, often termed "world simulators" due to their emergent understanding of physical dynamics and temporal coherence. Our work builds upon CogVideoX [48], a transformer-based latent diffusion model that employs a 3D Variational Autoencoder to compress videos across both spatial and temporal dimensions, enhancing coherence. To improve text-video alignment, CogVideoX also integrates an expert transformer with adaptive LayerNorm, enabling deep fusion of visual and textual modalities.
@@ 3.1 Temporal Editing Captions
To automate this process, we leverage recent advances in Vision-Language Models (VLMs) [10, 46, 45, 14, 38]. The VLM, specifically ChatGPT-4o [33], is instructed to produce a concise video scenario that highlights how elements within the image change or move over time. The generated caption captures the essential transformations while maintaining a static camera perspective unless movement is necessary. To improve generation quality, we utilize in-context learning (ICL) [12], providing the VLM with nine exemplar prompt-caption pairs. The complete prompt template, ICL examples, and an ablation study comparing this approach with directly using the target captions are included in Section B of the Appendix.
@@ 3.2 Video Generation
We employ CogVideoX (I2V-5B) [48], a pretrained generative video latent diffusion model utilizing a transformer-based architecture. Specifically, we use its image-to-video variant, which has been fine-tuned to generate videos starting from an input image $I_{s}$. During generation, $I_{s}$ is encoded and concatenated with noise in latent space, where the model applies a denoising process guided by the temporal caption $\tilde{c}$. As elaborated in Section 4, this conditioning allows generated videos to start from $I_{s}$ and evolve naturally along the image manifold, maintaining temporal coherence and consistency. Additionally, the model’s transformer architecture enables effective fusion of visual and textual information, allowing precise control over the editing process through our temporal captions. Formally, given the video generator $G$, we define the generation process as:
@@ 3.3 Frame Selection
To automate the selection of $t^{*}$ and avoid manual frame-by-frame review, we employ an automated approach. After generating the sequence $V$, we sample every fourth frame, imprinting each with a unique identifier and assembling them into an image collage alongside $I_{s}$. Inspired by [23], which introduces a novel approach to video comprehension by transforming videos into image grids, we use a VLM, specifically GPT-4o, to assist in selecting $t^{*}$ by providing it with the collage and the editing prompt $c$. The VLM is tasked with identifying the frame that best fulfills the editing intent, evaluating each frame’s alignment with $c$ and fidelity to $I_{s}$. The model is instructed to select the optimal frame with the lowest index that completes the edit. An ablation study, detailed in Section C of the appendix, evaluates the effectiveness of our automated approach against the baseline of selecting the final frame of the video.
Note that while we refer to an optimal $t^{*}$, the definition of the required edit can be subjective and dependent on user preferences. For example, in Figure 3, different users may select frames of the cat jumping based on variations in its altitude. Thus, frame selection could serve as a flexible and customizable advantage in certain scenarios.
@@ 4 Editing Manifold Pathway
We manually filter the outcomes to ensure alignment with descriptions, with examples provided in the supplementary material in Section D. Then, we use a CLIP ViT-B/32 model [36] to extract a 512-dimensional feature vector for each image. Using these features, along with 25 feature vectors extracted from random noise images, we perform Principal Component Analysis (PCA) to reduce the dimensionality into a two-dimensional subspace. The resulting subspace, now suitable for visualization, is depicted in Figure 2.
@@
Consider an original image from the AI group that we wish to edit using the target prompt: “A happy person making a heart shape with their hands”. Current editing methods generate a single image, which may cause an abrupt transition to the red cluster, effectively removing the ’AI’ on the shirt. This behavior is illustrated in the figure, where the edit by LEDITS++ is positioned near the red cluster, and the ’AI’ text on the shirt disappears. In contrast, our method leverages video generation to perform the edit smoothly, moving along the manifold with incremental changes until reaching the required edit in the purple cluster, as indicated by black arrows in Figure 2. The temporal editing caption guiding this process is: “A happy person very slowly raising their hands to form a heart shape”. This gradual progression lets us reach the purple cluster, resulting in an edited image where the person maintains the ‘AI’ on their shirt while making a heart shape—a faithful preservation of the original image’s key attributes. This experiment illustrates that the proposed paradigm enables smooth traversal across the image manifold, enabling consistent edits while preserving the essential characteristics of the original image.
@@ 5 Experiments
We evaluate Frame2Frame against state-of-the-art image editing methods, including LEdits++ [4], SDEdit [31], Pix2Pix-Zero [35] Imagic [22], and FLUX-based FlowEdit [24]. Our evaluation spans two benchmarks: the established TEdBench [22] for general image editing, and our newly introduced PosEdit, specifically designed for human pose editing. Finally, we conduct a human evaluation to assess our method based on real user preferences.
@@ 5.1 Evaluation Protocol
Table 2: PosEdit Results. Quantitative evaluation on PosEdit. Source metrics assess similarity to the original image, while Target metrics include LPIPS and CLIP-I comparisons to the ground-truth target image, along with the CLIP score for edit accuracy. “GT” provides ground-truth target image metrics for context.
@@ 5.2 TEdBench Evaluation Results
The qualitative advantages of our method are visually evident in the comparisons presented in Figure 4. The figure demonstrates our method’s superior performance across diverse editing scenarios, producing results that both faithfully execute the intended edits while maintaining strong alignment with the source image. Additional examples showcasing these capabilities are provided in the appendix in Section A.
@@ 5.3 PosEdit Benchmark
The proposed benchmark specifically focuses on human pose manipulation. Unlike TEdBench, each editing task in PosEdit includes a ground truth frame extracted from the same subject in the target pose. The availability of reference images allows for a more comprehensive evaluation of both editing accuracy and identity preservation. In the dataset, each editing task consists of two images: a source image showing the subject in a neutral standing pose with arms relaxed at their sides, and a ground-truth edited target image capturing the subject performing in a specific pose. The benchmark also provides a prompt for each task that describes the target pose the subject should achieve (e.g., "A person in a basketball shooting posture.").
Figure 5: Qualitative Results on PosEdit. Comparison between our Frame2Frame method and LEDITS++ on human motion editing tasks. For each example, we show the source image, edited results from both methods, and the ground-truth target image. Our method better preserves subject identity while achieving more natural pose transitions. The evaluation metrics for each image are provided in Section F of the appendix.
@@
The evaluation on PosEdit follows the same protocol established for TEdBench: generating multiple variants using nine different seeds and manually selecting the best result for each method. However, PosEdit’s ground-truth target images enable additional evaluation metrics. Beyond measuring preservation of source content (via LPIPS and CLIP-I between source and edited images) and edit accuracy (via CLIP score with the prompt), we compute LPIPS and CLIP-I metrics between the edited image and its corresponding ground-truth target. This enhanced evaluation directly assesses both the accuracy of the transformation and preservation of identity features. Additionally, we include the evaluation metrics of the ground-truth images as a reference to the rest of the results.
@@
Two important implementation notes: First, for Pix2Pix-Zero, which requires source image descriptions, we use the same static prompt across all tasks: "A person standing naturally with his arms relaxed at his sides." Second, we exclude Imagic [22] from this comparison as their official implementation is not publicly available and their results are only reported on TEdBench.
5.4 Human Evaluation Survey
To further evaluate our method, we conducted a human survey. As indicated in Table 1, LEDITS++ emerges as the most competitive method compared to ours, making it a natural choice for comparison. Each participant was presented with 20 randomly sampled TEdBench samples, including the source image, the target edit prompt, and the corresponding edited outputs from both methods. Inspired by the methodology proposed by [49], participants evaluated each comparison based on two criteria: (1) edit accuracy relative to the prompt, and (2) edit quality—defined as the preservation of visual fidelity to the source image, seamless integration of edited elements, and the overall natural appearance of modifications. We collected responses from 59 randomly selected online evaluators. To ensure an unbiased assessment, evaluators were not informed of the study’s objectives. The results were quantified using two metrics: (i) overall global preference, expressed as a percentage, and (ii) aggregated per-image preference, where a tie resulted in each method receiving 0.5 points.
@@ 5.4 Human Evaluation Survey
Table 3: Human Survey Results. Human evaluation on TEdBench shows that F2F surpasses LEDITS++ in edit accuracy while offering a significant advantage in preserving the original image.
Our results, summarized in Table 3, demonstrate that F2F outperforms LEdits++ on both metrics. Specifically: For edit quality, F2F achieved a global preference score of 53%, slightly surpassing the 47% obtained by LEdits++. In terms of edit accuracy, F2F achieved a 65.6% global preference compared to LEdits++’s 34.4%. The per-image preference results mirrored this trend, indicating robustness across individual examples and no significant influence of outliers. These findings reinforce our claim that smooth temporal editing—using video as a medium—preserves essential scene characteristics while successfully performing the edit. Our results suggest that the gap between F2F and LEdits++ in source preservation is larger than reflected by the LPIPS scores in Table 1. The full survey details are provided in the supplementary materials in Section G.
@@ 5.5 Additional Vision Tasks
We demonstrate our framework’s applicability beyond traditional editing by applying it to fundamental image manipulation tasks: denoising, deblurring, outpainting, and relighting. For these experiments, we utilize Runway Gen-3 [13] as our video generation backbone, as it showed superior performance on these specific tasks compared to other generative video models. Since these tasks serve as extensions and potential applications of our method, we focus on qualitative results, leaving quantitative evaluation for future work. Our results, shown in Figure 6, demonstrate strong performance across all tasks. We associate this success to the natural alignment between these operations and common video scenarios: deblurring maps to camera focusing, outpainting to camera motion, and relighting to time-lapse lighting changes. This alignment enables our method to leverage the video model’s learned temporal dynamics, achieving high-quality results without task-specific training. Full prompts are provided in Section I of the supplementary material.
Extended Results. Additional examples illustrating our method’s editing performance across diverse backgrounds are provided in Section H of the Appendix.